The difference of two numbers is 15. The first number is 5 more than 2 times the second. What are the two numbers?
Answer by ptaylor(2198)

Let x=second number
2x=2 times the second number
2x+5=5 more than 2 times the second
Then 2x+5=the first number
Now we are told that:
2x+5-x=15 subtract 5 from both sides
2x+5-5-x=15-5 collect like terms
x=10 -------------------------------second number
2x+5=2*10+5=25 ---------------------first number
CK
25-10=15
15=15

Hope this helps ----ptaylor